A population consists of three numbers (3, 6 ,9). Consider all possible samples of sizes 2 which can be drawn without replacement from the population. Find the following


1
Expert's answer
2021-10-18T17:15:08-0400

I.


"mean=\mu=\dfrac{3+6+9}{3}=6"

"variance=\sigma^2=\dfrac{1}{3}((3-6)^2+(6-6)^2"

"+(9-6)^2)=6"

"\sigma=\sqrt{\sigma^2}=\sqrt{6}"

II. There are "\dbinom{3}{2}=3" samples of size two which can be drawn without replacement: 


"\begin{matrix}
 Sample & Sample\ mean \\
 (3,6) & 4.5 \\
 (3,9) & 6 \\
 (6,9) & 7.5 \\
\end{matrix}"

III.


"\begin{matrix}
 \bar{X} & P(\bar{X}) \\
 4.5 & 1/3 \\
 6 & 1/3 \\
 7.5 & 1/3 \\
\end{matrix}"

IV.


"\mu_{\bar{X}}=4.5(1/3)+6(1/3)+7.5(1/3)=6"


"\sum_i\bar{X}_i^2P(\bar{X_i})=4.5^2(1/3)+6^2(1/3)+7.5^2(1/3)""=37.5"

"\sigma_{\bar{X}}^2=\sum_i\bar{X}_i^2P(\bar{X_i})-\mu_{\bar{X}}^2=37.5-6^2=1.5"

"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{1.5}"

"\mu_{\bar{X}}=6, \sigma_{\bar{X}}=\sqrt{1.5}"

V.

The mean "\mu_{\bar{X}}" and standard deviation "\sigma_{\bar{X}}" of the sample mean "\bar{X}" satisfy


"\mu_{\bar{X}}=6=\mu,"

"\sigma_{\bar{X}}=\sqrt{1.5}=\dfrac{\sqrt{6}}{\sqrt{2}}\sqrt{\dfrac{3-2}{3-1}}=\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}"